# Vejde: A Framework for Inductive Deep Reinforcement Learning Based on Factor Graph Color Refinement

### v1

@@ 6 Related Work
There have been a number of previous works that combine methods of machine learning on graphs with decision problems, and Vejde piece-wise overlaps with several of these. We summarize the primary differences and overlapping features with a non-exhaustive selection of related works in Table 5.
@@
There exists a large body of work that falls into the category of relational reinforcement learning. The book by Van Otterlo (2009) covers much of research in the area prior to the 2010s, which is composed of both purely symbolic and non-neural machine learning methods for both value-based and policy-based relational reinforcement learning. The survey by Mohan et al. (2024) covers more recent works, from the more generic perspective of utilizing structure in RL, as well as some older works. They use the term “side information” to refer to various forms of structural knowledge provided to agents to improve learning, but which are not part of the typical MDP formalism. Side information used by various works include data abstractions, information about system dynamics and goal formulations, which typically improve data efficiency or policy generalization at the cost of requiring more prior knowledge about the problem or computational power. By their categorization, the only side information we use in Vejde is data abstraction, where we assume that there exists a data description language that can describe object classes and relations from a given problem domain. This is, in our opinion, a significantly lighter assumption than for instance Sharma et al. (2023) makes, which is full knowledge of the dynamic Bayesian network in order to construct the input graph for the policy.
@@
An alternate method of inductive policy generation which we consider interesting is decision list learning. The state is represented with symbolic logic, but rules for actions are represented explicitly in a human-readable manner rather than implicitly encoded as weights in a neural network. Fern et al. (2006) presents a symbolic method that searches for a policy through a combination of beam-searching and policy iteration. Neural methods for decision list rule generation that incorporate numeric optimization have also been proposed (Hazra & Raedt, 2023; Delfosse et al., 2023), which in some cases assign numeric weights to the rules in order to make the policy probabilistic. Representing rules in a human-readable manner has an immediate benefit of being able to interpret the policy, which is good for transparency, but requires a method for searching a potentially large space of possible rules. We also recognize the possibility of using a trained MPNN policy to extract explicit rules post-hoc.
@@
The topic of generalization in reinforcement learning for classes of problems has also been covered from the perspective of deep learning. In their survey of RL generalization, Kirk et al. (2023) separate learning problems into three categories: singleton problems where the training and test environments are the same; IID generalization problems, where the test environments are different but sampled from the same distribution as the train environments; Out of Distribution (OOD) generalization problems, where the test environments are sampled from a different distribution. An example of such problems is the collection “ProcGen”, which define problems using distributions over variables such that randomly generated problem instances can be sampled from them (Cobbe et al., 2020).
@@ 6.1 Deep Relational Reinforcement Learning
The works which we consider closest in type to ours are those which use graph-based state representations to predict action probabilities in a model-free learning context. The method we use to sample actions is influenced by Janisch (2024). They use an autoregressive decoding scheme to sample the components of actions, which allows for both nullary and higher-order actions. For actions with an arity greater than one, additional rounds of message-passing are executed after sampling an object, marking objects that have already been selected with an additional feature. Unlike us, they use simple graphs to represent the facts of the state, which limits the method to problems with binary relations. Agents are trained using policy gradient optimization in an unsupervised manner. They evaluate their method on implementations of SysAdmin, Blockworld and a modified version of the game Sokoban.
@@
Garg et al. (2019), like us, use reinforcement learning to train policies based on the RDDL domains SysAdmin, Game of Life and Academic Advising. These are all represented as a simple graph of binary relations between objects, with a single unary action. While this work use a similar evaluation as us, we can not compare our results directly with them, as they too present normalized scores but have not calculated them in relation to any baseline policies. Therefore, we can not determine if their method performs better than a random or null policy on these domains. Additionally, they train and test on hand-picked subsets of instances, rather than all instances.
@@
Ammanabrolu & Riedl (2019) offers a somewhat different perspective, coming from a natural language processing context. Agents are evaluated on a text-based game, where an agent has to navigate a simulated space and solve puzzles. They represent observations using a simple graph, which use word embeddings as node features, and select actions as Resource Description Framework (RDF) triplets, on the form (subject, verb, object).
@@
We also wish to highlight a set of works we refer to as object-oriented reinforcement learning, which includes works by Guestrin et al. (2003)Diuk et al. (2008) and more recently Zambaldi et al. (2019). We categorize these works by the fact that they model the state using conditionally independent objects, but without explicit relationships. Mohan et al. (2024) classifies these works as factored representations, as opposed to relational representations that include relations, as we do. A number of works in the context of multi-agent reinforcement learning, such as Foerster et al. (2018), arguably fit this category when the complete system state is factorized into discrete states or observations for each individual agent. The object-oriented approach is practical in the sense that one does not have to define or observe relationships between objects, but tends to have worse scaling properties since approaches typically compares every entity to one another. It is, however, less sensitive to violations of the homophily assumptions which MPNNs rely on, i.e. that things which are connected in the graph are related.
@@ 6.2 Automated Planning
Many works that use symbolic logic to represent problems operate in the research area of planning, which tends to assume deterministic system models, as opposed to MDPs with probabilistic state transitions. Nevertheless, we share features with multiple works within this research area. For instance, Chen & Thiébaux (2024) and Ståhlberg et al. (2023) both use factor graph state representations, but predict heuristic scores over subsequent states rather than probabilities over actions. Chen et al. (2024) use the WL algorithm to generate latent states, which are then used to produce heuristic scores. In a subsequent work, Chen & Thiébaux (2024) extends the WL algorithm to incorporate numeric features, and evaluate their method on a single deterministic goal-oriented problem, Blockworld. They train their models in a supervised manner based on pre-calculated optimal plans. We find this pair of works interesting in that they show that a computationally simpler representation can be used to generate latent states, at least for the singular problem they investigated. In Ståhlberg et al. (2023) agents are trained in an unsupervised manner using reinforcement learning, as well as a value function using supervised learning. They evaluate their method on several problems defined in PDDL. In a subsequent work, Ståhlberg et al. (2025) focuses on methods for higher-order action selection, but these are only evaluated with supervised learning.
@@
Ammanabrolu & Riedl (2019) No No No Yes 2
Garg et al. (2019) No Yes No Yes 1
Ståhlberg et al. (2023) No No No No N/A
Chen & Thiébaux (2024) Yes No Yes No N/A
Janisch (2024) No Yes No Yes [0, 1, 2]
Ståhlberg et al. (2025) Yes No No No 2
@@ 7.1.1 Vejde
The Vejde agent trained with RL received average scores higher than the trivial policies on $\nicefrac{{8}}{{10}}$ domains. We consider this a positive indication that the method can be used for developing policies generalize to unseen problem instances within a class of problems. We observed a negative difference in score between the mimic policy and RL policy for most problem domains. Elevators has the most significant difference, with an 87 percent unit drop between the experiments. Since the problems and Vejde architectures were the same for both experiments, we attribute this difference to PPO not discovering a set of policy parameters that work and those found when imitating Prost. It remains an open question as to how this difference can be reduced. Navigation and CrossingTraffic occur in the set of problems the Vejde policies receive the lowest scores in, both in the reinforcement and imitation learning experiments. They are grid-based problems, with discrete cartesian coordinates represented as objects and only nullary actions. These are not the kinds of problems that we otherwise would choose to use Vejde with, in that representing a uniform grid as a graph is needlessly complicated and the number of possible actions is constant regardless of the grid size. Nevertheless, they fit the selection criteria we set, so we feel that it would be unfairly selective to exclude them. They are arguably a good test for the ability of the architecture to generate graph embeddings, in that they are essentially graph-classification tasks, but the hypothesis that a whole graph aggregation is needed to solve these problems remains untested.
@@ 7.1.2 MLP
The transductive MLP policies received higher average scores than Vejde agents for individual instances, which means that they represent a good option if structural generalization is not a priority. If the number of entities in the problem changes, more models need to be trained to account for the different input space, [^15] The practical benefits of having to handle one agent per domain as opposed to ten is hard to quantify, but not insignificant. or an input space that is agnostic to the number of entities need to be used. In theory, a Vejde policy can only be as good as a MLP policy on a given instance, assuming that the MLP policy is optimal, and the inductive policy is also optimal for all problem instances. If the inductive policy can not be optimal for all instances due to problem complexity, we would expect a score of the inductive policy to occur between those for the optimal policy for the instance and the trivial policies. We observe, on average, that the scores of the Vejde agents were statistically close to the scores of the MLP agents. In practice, due to the somewhat unstable optimization that deep RL constitutes, both the MLP and GNN may converge to local minima. This is our primary explanation as to why Vejde policies scored higher than MLP policies on some domains. We could strengthen our evaluation by training five copies of MLP policies with some variation, such as different initialization values, to pair with the five Vejde agents using different training sets. However, given the long training time required to train the 90 MLP policies, we opted against this. We hypothesize that our primary takeaway, that the scores of the inductive Vejde policies is correlated with the transductive MLP policies, will still hold.
@@ 7.1.3 Prost
Prost receives the highest scores on most problems, showing its strong capability to search for optimal actions. However, the main drawback of Prost is that it is always dependent on an accurate generative model of the problem during test-time inference. Depending on the use case, this can be difficult to accomplish due to not having an accurate model available for the problem instance, or that the execution environment does not have the resources to perform continuous simulations. By comparison, the RL policies only need the simulator during training but do not perform any planning when choosing actions. The cost of the high scores with Prost is also paid with a much higher inference time, which we had to count in days, as opposed to minutes for the RL policies. Despite its strong overall performance, Prost performs little better than the do-nothing policy on certain problem instances, most notably from the domain AcademicAdvising. We are unsure if this is caused by Prost not being well-adapted for this particular problem, or the instances being defined such that doing nothing is indeed the optimal course of action. The higher average scores of the Vejde agents suggests the latter may not be the case, however. The low scores of Prost on this domain is also reflected in the mimic policy, which are significantly worse on the problem compared to its RL counterpart.
7.2 Structured State Spaces
@@
The maximum arity of facts across the problems we evaluate with is two. We could therefore represent the state for all included problems as a simple graph consisting of only nodes and edges, as Janisch (2024) or Garg et al. (2019). This is a format which is also known as Knowledge graphs, or RDF graphs, within certain research contexts (Ammanabrolu & Riedl, 2019). It is not common for works that handle KGs to incorporate numeric features, or object-specific features at all, which we assume to have available. Fey et al. (2024a) presents a graph representation directly influenced by relational databases, which extends simple graphs to include heterogeneous data. They do, however, represent unary attributes as fixed-sized vectors, analogous to the table columns in relational databases. The bipartite representation we use is flexible in that entities can be represented with variably-sized sets of unary attributes. Being able to ignore missing attributes without padding is useful for problems where there are many possible object attributes, but only a small subset are present at a given time. We may, for instance, want to define many possible alerts for entities in a network intrusion detection system, where only the observed alerts are included in the agent input.
@@
In addition to changing the graph structure, there are also alternate methods for encoding the information in the graph which we considered. The conceptually simpler WL algorithm can be used instead of neural message-passing to encode the nodes and graph into latent states. However, for domains with continuous state variables the number of possible states becomes infinite, and we lose the ability to interpolate between latent states. Modifications can be made to the WL to allow for continuous representations (Chen & Thiébaux, 2024). We have also noted methods for belief propagation on factor graphs which incorporate neural networks, such as the one by Kuck et al. (2020), which could provide a different method for message-passing compared to the WL algorithm.
There are methods for generating object factored embeddings which eschew message-passing altogether. Deep sets (Zaheer et al., 2017) and Transformers (Vaswani et al., 2017) are both methods based on neural networks that can be used for factored representations without explicit relations. Both methods have the benefit that information from the entire state can be incorporated in the action selection, which also means that nullary facts can be incorporated without additional steps. However, they also possess worse scaling properties than MPNN, since every entity is compared with the others.
7.3 Problem Selection & Extensions
@@
In selecting the domains for evaluation, we excluded problems based on a set of filters, which we comment on in this subsection as stepping stones for extensions to the work. The problems defined in RDDL constitute an ideal situation for structural generalization. There are two main points which constitute this ideal: The first is that the dynamics and reward are specified in a factored and lifted manner, meaning that we can map objects to their respective types and not lose instance-specific information. The dynamics are also constant across problem instances. The second point is that the observations use the exact same predicates that are given to the agent, meaning that the value function and optimal policy can be defined using the language. In a more realistic setting, there may be a disconnect between the data abstractions and the underlying system. This arguably forms a partially observable MDP, or a “context” MDP as described by Kirk et al. (2023) which each require more complex solutions methods than for the fully observable problems we consider in this evaluation.
7.3.1 Continuous Actions
7.3.2 Higher-Order Actions
We define higher-order actions as actions which involve more than one object. If we set the limit to at most binary actions, predicting the action is roughly equivalent to link prediction in the greater context of graph learning. Implementations of higher-order actions tend to require that tuples of objects are compared, which scales poorly if done naively (Ståhlberg et al., 2025; Ammanabrolu & Riedl, 2019; Morris et al., 2019). For binary actions, the number of argument combinations scales quadratically, trinary cubically and so on. Autoregressive sampling of action arguments, as done by Janisch (2024), is an alternative method for producing higher-order actions which scales better than evaluating every possible tuple, but requires a solution for representing partial actions to the decoding model.
7.3.3 Concurrent Actions
The decentralized nature of MPNNs makes them suitable for implementing concurrent action selection, i.e. agents performing more than one action in a given timestep. Janisch (2024) demonstrated this with the SysAdmin domain, where the problem is changed to the agent selecting a subset of hosts rather than a single hosts. Having the agent select one action for each object, or a subset of objects, moves us closer to a cooperative multi-agent reinforcement learning formulation (Foerster et al., 2018), where we can view the graph as consisting of an arbitrary number of conditionally independent policies that should act in a cooperative manner.
7.3.4 Partial Observability
Including problems with partial observability, or hidden context variables as defined by Kirk et al. (2023) would, in our opinion, improve the practical usability of Vejde. Finding inductive solutions to problems with partial observability would likely require the addition of memory to the agent, or a temporal component to the database.
7.3.5 Action Preconditions
7.3.6 Nullary Domains
7.4 Problems not Defined in RDDL
While we do not focus on problems not defined in RDDL in this paper, our implementation is designed to be used for other, yet relational, problem areas. We chose to use RDDL for our evaluation as it allows us to evaluate Vejde on a varied selection of abstracted problem types without much additional implementation work. For the purpose of developing policies for automated defense, incident response has been simulated in a number of works, often as a two-agent game between an attacking and defending agent (Wolk et al., 2022). One of the more prominent examples of network incident response simulation is the CAGE set of problems, of which CAGE 4 is the most recent instance (Kiely et al., 2025). The problem state in both simulated and real-world incident response is often derived from log and system data, which tends to be relational by design and thus fits the relational reinforcement learning paradigm fairly well, as demonstrated by Thompson et al. (2024) for instance.

### v2

@@ 6 Related Work
There have been a number of previous works that combine machine learning on graphs with decision problems, and Vejde piece-wise overlaps with several of these. We summarize the primary differences and overlapping features with a non-exhaustive selection of related works in Table 5.
@@
There exists a large body of work in the realm of relational reinforcement learning. The book by vanotterloLogicAdaptiveBehavior2009a covers much of research in the area prior to the 2010s, which is composed of both purely symbolic and non-neural machine learning methods for both value-based and policy-based relational reinforcement learning. The survey by 10.1613/jair.1.15703 covers more recent works, as well as some older, from a more general perspective of incorporating structure in RL. They use the term “side information” to refer to various forms of structural information provided to agents to improve learning, but which is not part of the typical MDP formalism. Side information used by various works include data abstractions, information about system dynamics and goal formulations, which typically improve data efficiency or policy generalization at the cost of requiring more prior knowledge about the problem or computational power. By their categorization, the only side information we use in Vejde is data abstraction, where we assume that there exists a data description language that can describe object classes and relations from a given problem domain. This is, in our opinion, a significantly lighter assumption than for instance DBLP:conf/uai/SharmaAMS23 makes, which is full knowledge of the dynamic Bayesian network in order to construct the input graph for the policy.
@@
An alternate method of inductive policy generation which we consider interesting is decision list learning. The state is represented with symbolic logic, but rules for action selection are represented explicitly in a human-readable manner rather than implicitly encoded as weights in a neural network. DBLP:journals/jair/FernYG06 presents a symbolic method that searches for a policy through a combination of beam-searching and policy iteration. Neural methods for decision list rule generation that incorporate numeric optimization have also been proposed (DBLP:conf/pkdd/HazraR23; DBLP:conf/nips/DelfosseSDK23), which in some cases assign numeric weights to the rules in order to make the policy probabilistic. Representing rules in a human-readable manner has an immediate benefit of being able to manually analyze and interpret the policy, which is good for transparency, but requires a method for searching a potentially large space of possible rules. We also recognize the possibility to extract explicit rules from a trained MPNN policy post-hoc.
@@
The topic of generalization in reinforcement learning for classes of problems has also been covered from the perspective of deep learning. In their survey of RL generalization, DBLP:journals/jair/KirkZGR23 separate learning problems into three categories: singleton problems where the training and test environments are the same; Independent and Identically Distributed (IID) generalization problems, where the test environments are different but sampled from the same distribution as the train environments; Out of Distribution (OOD) generalization problems, where the test environments are sampled from a different distribution. An example of such problems is the collection “ProcGen”, which define problems using distributions over variables such that randomly generated problem instances can be sampled from them (DBLP:conf/icml/CobbeHHS20). By their classification, we are focusing on IID generalization of the agent in our evaluation.
@@ 6.1 Deep Relational Reinforcement Learning
The works which we consider closest in type to ours are those which use graph-based state representations to predict action probabilities in a model-free learning context. The sequential method we use to sample actions is influenced by janisch2024applications, who presents an autoregressive decoding scheme to sample the components of actions, which allows for both nullary and higher-order actions. For actions with an arity greater than one, additional rounds of message-passing are executed after sampling an object, marking objects that have already been selected with an additional feature. Unlike us, they use simple graphs to represent the facts of the state, which limits the method to problems with binary relations. Agents are trained using policy gradient optimization in an unsupervised manner. They evaluate their method on implementations of SysAdmin, Blockworld and a modified version of the game Sokoban.
@@
DBLP:conf/aips/GargBM19, like us, use reinforcement learning to train policies based on the RDDL domains SysAdmin, Game of Life and Academic Advising. These are all represented as a simple graph of binary relations between objects, with a single unary action. While this work use a similar evaluation as us, we can not compare our results directly with them, as they too present normalized scores but have not calculated them in relation to any baseline policies. Therefore, we can not determine if their method performs better than a random or null policy on these domains. Additionally, they train and test on hand-picked subsets of instances, rather than all instances.
@@
DBLP:conf/naacl/AmmanabroluR19 offers a somewhat different perspective, coming from a natural language processing context. Agents are evaluated on a text-based game, where an agent has to navigate a simulated space and solve puzzles. They represent observations using a simple graph, which use word embeddings as node features, and select actions as Resource Description Framework (RDF) triplets, on the form (subject, verb, object).
@@
We also wish to highlight a set of works we refer to as object-oriented reinforcement learning, which includes works by DBLP:conf/ijcai/GuestrinKGK03DBLP:conf/icml/DiukCL08 and more recently zambaldi2018deep. We categorize these works by the fact that they model the state using conditionally independent objects, but without explicit relationships. 10.1613/jair.1.15703 classifies these works as factored representations, as opposed to relational representations that include relations, as we do. A number of works in the context of multi-agent reinforcement learning, such as 10.5555/3504035.3504398, arguably fit this category when the complete system state is factorized into discrete states or observations for each individual agent. The object-oriented approach is practical in the sense that one does not have to define or observe relationships between objects, but tends to have worse scaling properties since approaches typically compares every entity to one another. It is, however, less sensitive to violations of the homophily assumptions which MPNNs rely on, i.e. that things which are connected in the graph are related.
@@ 6.2 Automated Planning
Many works that use symbolic logic to represent problems operate in the research area of planning, which tends to assume deterministic system models, as opposed to MDPs with probabilistic state transitions. Nevertheless, we share features with multiple works within this research area. For instance, DBLP:conf/nips/ChenT24 and DBLP:conf/kr/StahlbergBG23 both use factor graph state representations, but predict heuristic scores over subsequent states rather than probabilities over actions. DBLP:conf/icaps/ChenTT24 use the WL algorithm to generate latent states, which are then used to produce heuristic scores. In a subsequent work, DBLP:conf/nips/ChenT24 extends the WL algorithm to incorporate numeric features, and evaluate their method on a single deterministic goal-oriented problem, Blockworld. They train their models in a supervised manner based on pre-calculated optimal plans. We find this pair of works interesting in that they show that a computationally simpler representation can be used to generate latent states, at least for the singular problem they investigated. In DBLP:conf/kr/StahlbergBG23 agents are trained in an unsupervised manner using reinforcement learning, as well as a value function using supervised learning. They evaluate their method on several problems defined in PDDL. In a subsequent work, staahlberg2024learning focuses on methods for higher-order action selection, but these are only evaluated with supervised learning.
@@
DBLP:conf/naacl/AmmanabroluR19 No No No Yes 2
DBLP:conf/aips/GargBM19 No Yes No Yes 1
DBLP:conf/kr/StahlbergBG23 No No No No N/A
DBLP:conf/nips/ChenT24 Yes No Yes No N/A
janisch2024applications No Yes No Yes [0, 1, 2]
staahlberg2024learning Yes No No No 2
@@ 7.1.1 Vejde
The Vejde agent trained with RL received average scores higher than the trivial policies on $\nicefrac{{8}}{{10}}$ domains. We consider this a positive indication that Vejde can be used to represent policies that can generalize to unseen problem instances within a class of problems. We observed a negative difference in score between the mimic policy and RL policy for most problem domains. Elevators has the most significant difference, with an 87 percent unit drop between the experiments. Since the problems and Vejde architectures were the same for both experiments, we attribute this difference to PPO not discovering a set of policy parameters that work and those found when imitating Prost. It remains an open question as to how this difference can be reduced. Navigation and CrossingTraffic occur in the set of problems the Vejde policies receive the lowest scores in, both in the reinforcement and imitation learning experiments. They are both grid-based, with discrete cartesian coordinates represented as objects and only nullary actions. These are arguably not the kinds of problems that we would choose to use Vejde with, in that representing a uniform grid as a graph is needlessly complicated and the number of possible actions is constant regardless of the grid size. Nevertheless, they fit the selection criteria we set, so we feel that it would be unfairly selective to exclude them. They are arguably a good test of the architectures ability to generate whole-graph embeddings, in that they can be regarded as graph classification tasks, but the hypothesis that a whole graph aggregation is needed to solve these problems remains untested.
@@ 7.1.2 MLP
The transductive MLP policies received higher average scores than Vejde agents for individual instances, meaning that they represent a good alternative to Vejde if generalization is not a priority. If the number of entities in the problem changes, additional policies need to be trained to account for the different input space, [^17] The practical benefits of having to handle one agent per domain as opposed to ten is hard to quantify, but not insignificant. or an input space that is agnostic to the number of entities need to be used. In theory, a Vejde policy can only be as good as a MLP policy on a given instance, assuming that the MLP policy is optimal, and the inductive policy is also optimal for all problem instances. If the inductive policy can not be optimal for all instances due to problem complexity, we would expect a score of the inductive policy to occur between those for the optimal policy for the instance and the trivial policies. We observe, on average, that the scores of the Vejde agents were statistically close to the scores of the MLP agents. In practice, due to the somewhat unstable optimization that deep RL constitutes, both the MLP and GNN may converge to local minima. This is our primary explanation as to why Vejde policies scored higher than MLP policies on some domains. We could strengthen our evaluation by training five copies of MLP policies with some variation, such as different initialization values, to pair with the five Vejde agents using different training sets. However, given the long training time required to train the 90 MLP policies, we opted against this. We hypothesize that our primary takeaway, that the scores of the inductive Vejde policies is correlated with the transductive MLP policies, will still hold.
@@ 7.1.3 Prost
Prost receives the highest scores on most problems, showing its strong capability to search for optimal actions. However, the main drawback of Prost is that it is always dependent on an accurate generative model of the problem during test-time inference. This can be difficult to accommodate for use-cases where an accurate system model is not available, or the execution environment does not have the resources to perform continuous simulations. In contrast, the RL policies only need the simulator during training but do not perform any planning when choosing actions. The cost of the high scores with Prost is also paid with a much higher inference time, which we had to count in days, as opposed to minutes for the RL policies. Despite a strong overall performance, Prost performs little better than the do-nothing policy on certain problem instances, most notably from the domain AcademicAdvising. We are unsure if this is caused by Prost not being well-adapted for this particular problem, or the instances being defined such that doing nothing is indeed the optimal course of action. The higher average scores of the Vejde agents suggests the latter may not be the case, however. The low scores of Prost on this domain is also reflected in the mimic policy, which are significantly worse on the problem compared to its RL counterpart.
7.2 Alternative State Representations
@@
The maximum arity of facts across the problems we evaluate with is two. We could therefore represent the state for all included problems as a simple graph consisting of only nodes and edges, as janisch2024applications or DBLP:conf/aips/GargBM19. This is a format which is also known as Knowledge graphs, or RDF graphs, within certain research contexts (DBLP:conf/naacl/AmmanabroluR19). However, it is not common for works that handle KGs to incorporate numeric features, or object-specific features at all, which we require. Fey et al. (2024) presents a graph representation directly influenced by relational databases, which extends simple graphs to include heterogeneous data. They group unary attributes of entities into fixed-sized vectors, analogous to the table columns in relational databases, forming a heterogeneous graph. The bipartite representation we use is more flexible in that entities can be represented with variably-sized sets of unary attributes. Being able to ignore missing or unseen attributes without padding is practical for problems where there are many possible object attributes, but only a small subset are present at a given time. We may, for instance, want to define many possible alerts for hosts in a network intrusion detection system, while only including the alerts that have actually been observed in the agent input.
@@
In addition to changing the graph structure, there are also alternate methods for encoding the information in the graph, such as the graph attention method we describe in Appendix H. The conceptually simpler WL algorithm can be used instead of neural message-passing to encode the nodes and graph into latent states. However, for domains with continuous state variables the number of possible states becomes infinite, and we lose the ability to interpolate between latent states. Modifications can be made to the WL to allow for continuous representations (DBLP:conf/nips/ChenT24). We have also noted methods for belief propagation on factor graphs which incorporate neural networks, such as the one by Kuck et al. (2020), which could provide a different method for message-passing compared to the WL algorithm.
7.3 Limits to Generalization
@@
Problems defined in RDDL arguably represent an ideal situation for structural generalization. The dynamics and reward are specified in a factored and lifted manner, meaning that we can map objects to their respective types and not lose instance-specific information. The dynamics are also constant across problem instances and the observations use the exact same predicates that are given to the agent, meaning that the value function and optimal policy can be defined using the language. In a realistic setting, there may be a disconnect between the data abstractions and the system they represent. This arguably forms a partially observable MDP, or a “context” MDP as described by DBLP:journals/jair/KirkZGR23 which each require more complex solutions methods than for the fully observable problems we consider in this evaluation.